# Not Every Gift Comes in Gold Paper or with a Red Ribbon: Exploring Color Perception in Text-to-Image Models

### v1

@@ Appendix A Additional Details of our Benchmark
Appendix B Evaluation Metric
@@
After completing these steps, we compute three evaluation metrics for each color: CIELAB Distance($\Delta E_{CMC}$), RGB L2 Distance, and Accuracy (as defined in the main paper). Figure S3 illustrates the accuracy metric threshold. We filter images that do not accurately depict the two target objects. Specifically, we use SAM to perform object segmentation on the generated images with respect to the original prompt’s objects. We then filter images which do not have two object masks or masks with IOU larger than 0.5 to avoid cases of overlapping masks for the same object.
@@
Apart from the three main metrics reported, we evaluate some additional metrics detailed as follows.
@@ C.1.2 Editing Procedure
We use $\gamma_{AL}$ = 20 when $t<=35$ and $\lambda_{CL}$ = 112.5 when $t>=25$ otherwise 0. This means that the Stroop attention loss is only applied for the first 35 time steps, and the color loss is only applied in the later iterations after the first 25 steps. The rationale is that the cross-attention maps should align with the relevant object’s pseudo-ground-truth maps $A^{simp}_{object}$ early in the process; otherwise, the result may drift. We stop the Attention Loss after 35 steps to allow flexibility to the diffusion process and to allow the color loss to be more dominant. We start the color loss later because it depends on the interaction between the color and object cross-attention maps with respect to the relevant object on the image. Lastly, we perform blending at time $t_{b}$ to preserve background appearance as described in Equation 6 in the main paper.
@@ C.2.3 Post-Generation Editing Methods
In addition to the methods above, we further compare our editing method with AnySD [36], FPE [18] and MasaCtrl [6]. We used the default setup from the official GitHub repository for them. Additionally, for MasaCtrl, we followed their recommendation to inject the reference latents from the inversion process to the backward process for better reconstruction and editing results. To enable multi-object color editing with AnySD, we follow their official protocol by applying two consecutive single-object edits, one for each color-object pair. Runtime for all three models is roughly 25–30 seconds.
@@ Appendix D Additional Results
D.1 Color Leakage Experiment
@@ D.2 Evaluation Using Different Metrics
Method Ours(Imp $\uparrow$) AnySD (Imp $\uparrow$) FPE (Imp $\uparrow$) MasaCtrl (Imp $\uparrow$)
@@
SD 1.4 6.15\46.88\0.21 0.35\-0.51\0.03 1.42\8.26\0.03 0.36\0.21\0.01
SD 1.5 6.04\38.39\0.18 0.76\-1.89\-0.01 1.83\3.73\0.03 1.05\-0.82\-0.02
SD 2.1 7.43\51.02\0.22 1.63\3.48\0.03 2.12\7.94\0.03 0.57\1.82\-0.02
FLUX 4.11\30.12\0.16 1.46\2.06\0.04 1.72\7.54\0.07 1.28\3.61\0.02
A&E 6.40\33.16\0.15 0.47\-2.18\0.00 2.43\5.42\0.02 0.80\0.38 \-0.04
StructDiff 4.66\32.81\0.23 -0.18\-4.00\0.03 1.38\5.09\0.06 0.80\0.96 \0.02
SynGen 4.23\38.19\0.22 0.17\1.01\-0.04 2.89\18.83\0.16 1.49\7.37\0.08
RichText 7.91\29.59\0.23 2.88\9.38\0.11 5.06\11.72\0.15 3.24\5.77\0.07
BA 3.57\34.51\0.11 0.48\-1.13\-0.03 1.04\8.56\0.02 0.34\3.34\-0.01
@@
SD 1.4 9.52\70.14\0.20 0.72\4.58\0.03 0.32\0.075\-0.00 -0.17\-2.56\-0.03
SD 1.5 10.15\66.24\0.19 1.36\3.76\0.06 0.53\-1.14\0.00 0.35\-2.60\-0.02
SD 2.1 12.43\79.42\0.24 2.43\8.47\0.07 1.51\1.93\0.03 0.29\-0.56\0.00
FLUX 5.28\36.94\0.14 1.93\0.94\0.06 1.44\3.96\0.05 1.23\1.61\0.04
A&E 9.75\50.13 \0.23 0.82\-0.93\0.02 1.27\1.89\0.04 1.20\0.28\-0.01
StructDiff 7.98\44.88\0.18 -0.81\-7.70\-0.02 0.1\-2.43\-0.03 -0.60\-6.66 \-0.03
SynGen 3.05\26.78\0.17 -0.29\0.80\0.04 1.85\9.56\0.12 0.95\5.52\0.09
RichText 8.17\28.53\0.24 4.15\3.94\0.09 5.95\7.61\0.10 2.07\2.48\0.02
BA 8.03\61.89\0.18 2.91\13.29\0.06 0.43\-0.95\-0.01 0.032\-3.20 \-0.04
@@ D.2.1 LAB distance, RGB Mean and Accuracy
D.3 Evaluation Using Common Metrics in Image Generation
In addition to evaluating performance on the color-based metrics described in the main paper, we compare editing performance using CLIP Score (CS), CLIP Directional Score (CDS), and VQAScore improvement (edited compare to initial), see Table S3.
Our results show that while our CS and CDS scores are lower than FPE, this does not necessarily indicate inferior performance. As discussed in Figure 2 of the main paper, these metrics are often coarse and struggle to accurately capture fine-grained semantics, particularly in tasks involving precise color control. Thus, their lower scores may not be fully reflective of the effectiveness of our approach in this specific color-focused evaluation.
To address the limitations of these traditional metrics, we also incorporate VQAScore [17], which considered the SOTA for measuring similarity between image and text. Our editing method outperforms or is comparable to AnySD, FPE and MasaCtrl across all model datasets, except for FLUX-close, where the difference is minimal. However, even VQAScore, despite being a more advanced metric, still struggles to fully capture fine-grained semantic alignment, as illustrated in Figure 2 of the main paper. The limitations of existing evaluation metrics further emphasize the importance of our explicit evaluation pipeline, which directly quantifies perceptible semantic differences in color alignment, providing a more reliable assessment of model performance in compositional color tasks.
@@ D.3 Evaluation Using Common Metrics in Image Generation
SD 1.4 0.0035\0.010\0.08 0.0015\-0.0031\0.02 0.0036\0.018\0.01 0.0031\0.0080\0.01
SD 1.5 0.0036\0.0088\0.08 0.0021\0.0042\0.03 0.0031\0.019\0.01 0.0026\0.0049\0.00
SD 2.1 0.0029\0.016\0.10 -0.00063\0.0055\0.03 0.0038\0.017\0.02 0.0030\0.0044\-0.01
FLUX 0.0074\0.024\0.00 0.0021 \0.028\0.00 0.0049\0.026\0.01 -0.00081\0.015\-0.06
A&E 0.0032\0.014\0.10 0.0041\0.0058\0.03 0.0049\0.030\0.02 0.0022\0.014\0.00
StructDiff 0.0046\0.012\0.10 0.0069\0.0034\0.03 0.0038\0.022\0.02 0.0054\0.0088\0.01
SynGen 0.0054\0.026\0.03 0.00087\-0.0040\-0.01 0.0073\0.038\0.03 0.0063\0.017\0.01
RichText 0.013\0.051\0.12 0.018\0.014\0.07 0.016\0.062\0.07 0.011\0.029\0.00
BA 0.0025\0.0081\0.04 0.0023\0.00035\0.01 0.0034\0.022\0.01 -0.00025\0.0063\-0.03
@@
SD 1.4 0.00095\-0.0096\0.12 0.0015\-0.0020\0.03 0.0032\0.018\0.00 0.0020\0.0076\-0.01
SD 1.5 0.0013\-0.0022\0.13 0.0013\0.0046\0.03 0.0030\0.018\0.00 0.0021\0.0053\-0.01
SD 2.1 0.0019\0.00018\0.14 -0.00055\0.00020\0.04 0.0044\0.022\0.01 0.0020\0.010\-0.01
FLUX 0.0080\0.026\0.03 0.0050\0.029\0.01 0.0082\0.041\0.00 0.0051\0.023\-0.03
A&E 0.0024\0.014\0.13 0.0033\0.011\0.02 0.0051\0.030\0.01 0.0027\0.012\-0.01
StructDiff 0.0024\0.0084\0.16 0.0054\0.01\0.03 0.0040\0.027\0.02 0.0041\0.014\-0.01
SynGen 0.0075\0.027\0.00 0.0011\0.0030\0.00 0.0078\0.039\-0.02 0.0057\0.016\0.00
RichText 0.022\0.070\0.21 0.022\0.032\0.14 0.021\0.079\0.13 0.011\0.028\0.03
BA 0.000059\-0.0064\0.13 0.0024\0.0048\0.04 0.0035\0.016\0.00 0.0020\0.0059\-0.01
@@
D.4 Comparison with ColorPeel
@@ D.6 Performance without using SAM masks
D.7 Additional Qualitative Results
D.7.1 Propriety Models Experiment
OpenAI 4o Source Refer to caption    +Ours Refer to caption +AnySD Refer to caption +FPE Refer to caption +MasaCtrl Refer to caption Gemini2.5 Refer to caption    Refer to caption Refer to caption Refer to caption Refer to caption Grok3 Refer to caption    Refer to caption Refer to caption Refer to caption Refer to caption

### v2

@@ Appendix A Additional Details of our Benchmark
Appendix B Evaluation Protocol and Metrics
@@
After completing these steps, we compute three evaluation metrics for each color: CIELAB Distance($\Delta E_{CMC}$), RGB L2 Distance, and Accuracy (as defined in the main paper). Further details on threshold selection for the accuracy metric are provided in Section D.10.
We exclude source–edited image pairs if either image fails to accurately depict the two target objects. Specifically, we use SAM to perform object segmentation on the generated images with respect to the original prompt’s objects. We then filter images which do not have two object masks or masks with IOU larger than 0.5 to avoid cases of overlapping masks for the same object. This filtering may introduce minor variations in accuracy across models, but ensures a fair and reliable evaluation.
@@
Additional metrics. Beyond the primary fidelity metrics above, we also use:
@@
Image-Reward. We calculated ImageReward [39] of the edited image with respect to the original text prompt (”a color1 colored object1 and a color2 colored object2”). ImageReward is a learned reward model trained on 137k expert-labeled human preference comparisons, which assigns higher scores to images that are both visually pleasing and semantically aligned with their text prompts.
These metrics will be used throughout our experiments. Quantitative comparisons and analysis are reported in Section D.2.
@@ C.1.2 Editing Procedure
We use $\gamma_{AL}$ = 20 when $t<=35$ and $\lambda_{CL}$ = 112.5 when $t>=25$ otherwise 0. This means that the Stroop attention loss is only applied for the first 35 time steps, and the color loss is only applied in the later iterations after the first 25 steps. The rationale is that the cross-attention maps should align with the relevant object’s pseudo-ground-truth maps $A^{simp}_{object}$ early in the process; otherwise, the result may drift. We stop the Attention Loss after 35 steps to allow flexibility to the diffusion process and to allow the color loss to be more dominant. We start the color loss later because it depends on the interaction between the color and object cross-attention maps with respect to the relevant object on the image.
Finally, we want to ensure that our modifications do not alter the structure of the original image. To preserve the structural integrity of the original image, we draw inspiration from prior work in diffusion-based image editing [16, 7, 21] and replace the self-attention for the full prompt with the self-attention obtained using the simplified prompt. In addition, to maintain the background, we perform blending at time $t_{b}$ similar to [29] in order to preserve the background appearance during the editing process (see Figure 3 in the main paper). Concretely, for each object, we use the object masks $M^{simp}_{object}$ introduced previously in Section 4.2.2 in main paper and take the union to obtain $M^{simp}_{fg}$. Then, we perform blending at time $t_{b}$ during the backward process, updating the latent as follows:
$Z_{t_{b}}\rightarrow Z_{t_{b}}\cdot{M^{simp}_{fg}}+Z^{simp}_{t_{b}}\cdot({1-M^{simp}_{fg}})$ (S4)
@@ C.2.3 Post-Generation Editing Methods
In addition to the methods above, we further compare our editing method with AnySD [41], MagicBrush [42], InstructPix2Pix[5], FPE [21] and MasaCtrl [7]. We used the default setup from the official GitHub repository for them. Additionally, for MasaCtrl, we followed their recommendation to inject the reference latents from the inversion process to the backward process for better reconstruction and editing results. To enable multi-object color editing with AnySD. For AnySD and MagicBrush we followed the official instruction prompt from AnySD applying two consecutive single-object edits, one for each color-object pair. For InstructPix2Pix we follow their official editing prompt, applying two consecutive single-object edits, one for each color-object pair as it resulted in the best performance.
@@ Appendix D Additional Results
D.1 Analysis of Color Leakage
@@ D.2 Evaluation Using Different Metrics
Method Ours(Imp $\uparrow$) AnySD (Imp $\uparrow$) MagicBrush (Imp $\uparrow$) InstructPix2Pix (Imp $\uparrow$) FPE (Imp $\uparrow$) MasaCtrl (Imp $\uparrow$)
@@
SD 1.4 6.15\46.88\0.21 0.35\-0.51\0.03 3.35\18.72\0.10 0.79\11.94\0.00 1.42\8.26\0.03 0.36\0.21\0.01
SD 1.5 6.04\38.39\0.18 0.76\-1.89\-0.01 3.29\11.52\0.10 2.30\10.72\0.04 1.83\3.73\0.03 1.05\-0.82\-0.02
SD 2.1 7.43\51.02\0.22 1.63\3.48\0.03 3.54\17.59\0.11 3.74\21.84\0.04 2.12\7.94\0.03 0.57\1.82\-0.02
FLUX 4.11\30.12\0.16 1.46\2.06\0.04 0.97\5.01\0.02 -3.22\-5.62\-0.18 1.72\7.54\0.07 1.28\3.61\0.02
A&E 6.40\33.16\0.15 0.47\-2.18\0.00 5.00\14.62\0.11 2.30\3.56\-0.05 2.43\5.42\0.02 0.80\0.38 \-0.04
StructDiff 4.66\32.81\0.23 -0.18\-4.00\0.03 2.93\13.75\0.14 1.82\9.12\0.07 1.38\5.09\0.06 0.80\0.96 \0.02
SynGen 4.23\38.19\0.22 0.17\1.01\-0.04 1.94\13.07\0.10 0.51\15.80\-0.01 2.89\18.83\0.16 1.49\7.37\0.08
RichText 7.91\29.59\0.23 2.88\9.38\0.11 4.88\3.96\0.16 3.29\0.24\0.05 5.06\11.72\0.15 3.24\5.77\0.07
BA 3.57\34.51\0.11 0.48\-1.13\-0.03 1.47\12.77\0.01 0.29\5.65\-0.09 1.04\8.56\0.02 0.34\3.34\-0.01
@@
SD 1.4 9.52\70.14\0.20 0.72\4.58\0.03 3.49\18.90\0.15 -3.42\-9.41\-0.12 0.32\0.075\-0.00 -0.17\-2.56\-0.03
SD 1.5 10.15\66.24\0.19 1.36\3.76\0.06 4.52\27.08\0.14 -2.58\-10.92\-0.08 0.53\-1.14\0.00 0.35\-2.60\-0.02
SD 2.1 12.43\79.42\0.24 2.43\8.47\0.07 7.84\37.50\0.20 -2.56\-10.15\-0.06 1.51\1.93\0.03 0.29\-0.56\0.00
FLUX 5.28\36.94\0.14 1.93\0.94\0.06 -2.03\-12.73\-0.02 -13.18\-66.85\-0.29 1.44\3.96\0.05 1.23\1.61\0.04
A&E 9.75\50.13 \0.23 0.82\-0.93\0.02 4.65\5.64\0.13 -4.53\-31.02\-0.13 1.27\1.89\0.04 1.20\0.28\-0.01
StructDiff 7.98\44.88\0.18 -0.81\-7.70\-0.02 3.92\2.97\0.04 -4.80\-27.28\-0.16 0.1\-2.43\-0.03 -0.60\-6.66 \-0.03
SynGen 3.05\26.78\0.17 -0.29\0.80\0.04 -1.44\-1.80\0.01 -9.62\-38.78\-0.25 1.85\9.56\0.12 0.95\5.52\0.09
RichText 8.17\28.53\0.24 4.15\3.94\0.09 5.96\5.36\0.16 -1.66\-32.58\-0.04 5.95\7.61\0.10 2.07\2.48\0.02
BA 8.03\61.89\0.18 2.91\13.29\0.06 1.29\13.47\0.12 -9.09\-44.34\-0.15 0.43\-0.95\-0.01 0.032\-3.20 \-0.04
@@
Method Ours AnySD MagicBrush InstructPix2Pix FPE MasaCtrl
Close
SD 1.4 0.48 0.51 0.38 0.38 0.48 0.45
SD 1.5 0.63 0.67 0.41 0.50 0.62 0.53
SD 2.1 0.84 0.88 0.62 0.37 0.86 0.74
FLUX 1.26 1.38 0.93 0.98 1.39 0.91
A&E 0.48 0.51 0.14 0.11 0.50 0.37
StructDiff 0.53 0.59 0.43 0.45 0.56 0.52
SynGen 0.69 0.50 0.28 0.24 0.71 0.71
RichText 0.29 0.21 0.06 0.23 0.25 -0.00
BA 0.47 0.57 -0.57 0.16 0.59 0.44
Close-Mean 0.63 0.65 0.30 0.38 0.66 0.52
Distant
SD 1.4 0.60 0.66 0.13 0.15 0.61 0.54
SD 1.5 0.65 0.58 0.18 0.18 0.60 0.53
SD 2.1 0.79 0.78 0.41 0.40 0.79 0.61
FLUX 1.36 1.45 0.87 0.73 1.44 1.22
A&E 0.71 0.61 0.02 0.00 0.66 0.58
StructDiff 0.59 0.60 0.22 0.32 0.57 0.52
SynGen 0.92 0.80 0.27 0.06 0.98 0.95
RichText 0.38 0.15 -0.02 0.01 0.22 -0.15
BA 0.53 0.58 -0.83 -0.19 0.61 0.45
Distant-Mean 0.73 0.69 0.14 0.18 0.72 0.58
Table S3: Evaluation using image reward, comparing our method with AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl. We report the image reward mean score; higher is better.
In Table S3 we quantify image quality using Image Reward. Our evaluation shows that our model yields comparable performance while achieving significantly more accurate color edits. The largest margin is observed in distant-color prompts, where color edits are more challenging and prone to quality degradation. These results suggest that our edits do not harm—and may often enhance—visual fidelity.
In Table S4 we compare the editing performance of the different models using CLIP Score (CS), CLIP Directional Score (CDS), and VQAScore improvement (edited compare to initial), see Table S4. Our results show that while our CS and CDS scores are lower than FPE and InstructPix2Pix, this does not necessarily indicate inferior performance. As discussed in Figure 2 of the main paper, these metrics are often coarse and struggle to accurately capture fine-grained semantics, particularly in tasks involving precise color control. Thus, their lower scores may not be fully reflective of the effectiveness of our approach in this specific color-focused evaluation.
To address the limitations of these traditional metrics, we also incorporate VQAScore [20], which considered the SOTA for measuring similarity between image and text. Our editing method outperforms or is comparable to AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl across all model datasets, except for FLUX-close, where the difference is minimal. However, even VQAScore, despite being a more advanced metric, still struggles to fully capture fine-grained semantic alignment, as illustrated in Figure 2 of the main paper. The limitations of existing evaluation metrics further emphasize the importance of our explicit evaluation pipeline, which directly quantifies perceptible semantic differences in color alignment, providing a more reliable assessment of model performance in compositional color tasks.
@@
SD 1.4 0.0035\0.010\0.08 0.0015\-0.0031\0.02 -0.00097\0.0084\0.05 -0.00028\0.025\0.036 0.0036\0.018\0.01 0.0031\0.0080\0.01
SD 1.5 0.0036\0.0088\0.08 0.0021\0.0042\0.03 -0.0018\0.015\0.054 -0.0017\0.023\0.05 0.0031\0.019\0.01 0.0026\0.0049\0.00
SD 2.1 0.0029\0.016\0.10 -0.00063\0.0055\0.03 -0.0041\0.018\0.066 -0.0025\0.030\0.067 0.0038\0.017 \0.02 0.0030\0.0044\-0.01
FLUX 0.0074\0.024\0.00 0.0021 \0.028 \0.00 -0.0029\0.036 \-0.054 -0.0015\0.036 \-0.069 0.0049\0.026\0.01 -0.00081\0.015\-0.06
A&E 0.0032\0.014\0.10 0.0041\0.0058\0.03 -0.0011\0.024\0.035 -0.0028\0.033 \0.029 0.0049\0.030\0.02 0.0022\0.014\0.00
StructDiff 0.0046\0.012\0.10 0.0069\0.0034\0.03 0.0015\0.016\0.059 0.0064\0.024\0.049 0.0038\0.022\0.02 0.0054\0.0088\0.01
SynGen 0.0054\0.026\0.03 0.00087\-0.0040\-0.01 0.0043\0.020\-0.05 0.0013\0.023\-0.035 0.0073\0.038\0.03 0.0063\0.017\0.01
RichText 0.013\0.051\0.12 0.018\0.014\0.07 0.016\0.060\0.084 0.021\0.070 \0.093 0.016\0.062 \0.07 0.011\0.029\0.00
BA 0.0025\0.0081\0.04 0.0023\0.00035\0.01 -0.021\0.0080\-0.12 -0.0046\0.024 \-0.011 0.0034\0.022\0.01 -0.00025\0.0063\-0.03
@@
SD 1.4 0.00095\-0.0096\0.12 0.0015\-0.0020\0.03 -0.0061\0.0012\0.033 -0.0083\0.014\-0.04 0.0032\0.018\0.00 0.0020\0.0076\-0.01
SD 1.5 0.0013\-0.0022\0.13 0.0013\0.0046\0.03 -0.0038\0.015\0.062 -0.0045\0.026\-0.045 0.0030\0.018\0.00 0.0021\0.0053\-0.01
SD 2.1 0.0019\0.00018\0.14 -0.00055\0.00020\0.04 -0.0044\0.016\0.073 -0.0066\0.023 \-0.0025 0.0044\0.022\0.01 0.0020\0.010\-0.01
FLUX 0.0080\0.026\0.03 0.0050\0.029\0.01 0.00031\0.034\-0.11 -0.0012\0.041\-0.22 0.0082\0.041\0.00 0.0051\0.023\-0.03
A&E 0.0024\0.014\0.13 0.0033\0.011\0.02 -0.0058\0.028\-0.038 -0.0072\0.030\-0.11 0.0051\0.030\0.01 0.0027\0.012\-0.01
StructDiff 0.0024\0.0084\0.16 0.0054\0.01\0.03 0.0018\0.021\0.059 0.0034\0.029\0.0016 0.0040\0.027\0.02 0.0041\0.014\-0.01
SynGen 0.0075\0.027\0.00 0.0011\0.0030\0.00 0.00054\0.026\-0.14 -0.0041\0.019\-0.19 0.0078\0.039\-0.02 0.0057\0.016\0.00
RichText 0.022\0.070\0.21 0.022\0.032\0.14 0.030\0.093\0.11 0.031\0.097\0.027 0.021\0.079\0.13 0.011\0.028\0.03
BA 0.000059\-0.0064\0.13 0.0024\0.0048\0.04 -0.025\0.0031\-0.089 -0.011\0.0089\-0.13 0.0035\0.016\0.00 0.0020\0.0059\-0.01
@@
D.3 Comparison with ColorPeel
@@
D.4 Runtime Comparison
ColorEdit-Ours AnySD MagicBrush InstructPix2Pix FPE MasaCtrl
70 11 25 25 20 22
Table S5: Runtime comparison (in seconds) between our method and AnySD, MagicBrush, InstructPix2Pix, FPE and MasaCtrl.
In Table S5 we added a runtime comparison between our method and the baselines. As our method optimizes the latents for performing each color edit, it is indeed a bit slower than prior approaches. However, our approach delivers substantially higher color fidelity in comparison to prior work, remaining practical in terms of runtime for editing scenarios where quality is prioritized over speed.
@@ D.6 Performance without using SAM masks
D.7 Inversion-free Performance
Method Inversion(Imp $\uparrow$) w/o Inversion(Imp $\uparrow$)
Close
SD 1.5 6.04\38.39\0.18 3.75\37.07\0.21
Distant
SD 1.5 10.15\66.24\0.19 7.00\54.97\0.18
[…]
SD1.5 Refer to caption
+Ours w/o Inversion Refer to caption
”a pale-turquoise colored balloon and a hot-pink colored bench”
Figure S5: Editing results using our method without inversion. Images are generated with SD1.5, showing that our approach remains effective even without inversion.
In the previous section, we demonstrated that our method is robust to different inversion techniques. However, since inversion itself may occasionally fail, we further show that our approach does not depend on inversion at all. Specifically, we conduct an experiment where editing is performed without inversion: rather than inverting the image to obtain latents for editing, we generate an image from random noise using the input prompt and then apply our editing procedure on the same random initial latents. This experiment highlights the generality of our method, demonstrating its effectiveness even without inversion. We perform this evaluation using SD1.5, which serves as the base model for our approach. SAM-based segmentation is not employed in this setting, since image generation and editing occur simultaneously. As shown in Table S9, our method achieves strong performance without inversion, surpassing the standard SD1.5 baseline and confirming both the generality and inversion-independence of our color editing framework. Figure S5 shows that our editing method performs well without inversion, editing the right colors to the balloons and the bench.
D.8 Analysis of Overlapping Objects
Method Iou>> 0.5 0.1<< Iou $\leq$ 0.5 Iou $\leq$ 0.1
Close
[…]
Distant
[…]
Table S10: ColorEdit evaluation with respect to different overlap levels between the objects. We report the absolute improvement in Accuracy(ACC); higher is better, and 0 means no improvement. In brackets we report the relative amount of images with such overlap.
We added an analysis of overlapping objects, considering multiple overlap levels. In Table S10 the analysis shows that our model’s performance remains relatively stable even at high overlap levels ($IOU\geq 0.1$), still consistently improving performance over the source images. For Bounded-Attention we see low overlap as we used the default bounding rects from the official implementation which forces non overlapping objects.
D.9 Statistical Significance
We performed a paired non-parametric analysis, comparing images edited by our method with both the source images and those produced by baseline approaches (AnySD, MagicBrush, InstructPix2Pix, FPE, and MasaCtrl). The Wilcoxon signed-rank test was applied to paired LAB distance values under the same set of prompts. The results show that our method achieves statistically significant improvements over all baselines ($p<0.05$, Wilcoxon signed-rank test).
D.10 Accuracy Threshold Validation
Figure S6: Example user study question: participants were asked whether the difference between the reference object’s color and the color shown on the right was minor or clearly perceptible.
Metric Th-5 Th-10 Th-15 Th-20 Th-25
[…]
AUC 0.64 0.67 0.58 0.55 0.55
G-mean 0.53 0.66 0.54 0.5 0.46
Balanced Accuracy 0.64 0.67 0.58 0.55 0.55
Table S11: User study statistics in comparison with different thresholds, we see that threshold 10 is providing the best statistics in terms of AUC, G-mean and Balanced Accuracy.
To validate that our selected threshold aligns with human perception, we conducted a user study. Participants were asked whether an object’s color was sufficiently similar to the target text or clearly different, see Figure S6. In Table S11 we compared their responses against multiple thresholds and found that a threshold of 10 provides the best balance between the true positive and true negative rates and best overall performance in terms of AUC, G-mean, and balanced accuracy.
D.11 Additional Qualitative Results
D.11.1 Propriety Models Experiment
OpenAI 4o Source Refer to caption +Ours Refer to caption +AnySDRefer to caption +MagicBrushRefer to caption +InstPix2Pix Refer to caption +FPE Refer to caption
+MasaCtrl Refer to caption
Gemini2.5 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Grok3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption